### v1

@@ 3.6 Simulation Evidence
rd2d 1 17.497 -0.000 0.049 0.049 0.942 0.287
5 15.504 0.001 0.037 0.037 0.953 0.220
10 13.399 0.002 0.035 0.035 0.945 0.197
15 13.932 -0.000 0.031 0.031 0.947 0.175
21 15.490 0.002 0.043 0.043 0.962 0.292
25 13.964 -0.001 0.028 0.028 0.951 0.163
30 13.146 -0.000 0.033 0.033 0.959 0.190
35 14.536 0.002 0.036 0.036 0.954 0.207
40 16.961 0.000 0.046 0.046 0.949 0.264
Uniform 0.937 0.313
rd2d.dist kink = "off" 1 34.255 0.030 0.037 0.047 0.950 0.258
5 26.741 0.015 0.036 0.039 0.955 0.240
10 19.566 0.003 0.034 0.034 0.956 0.257
15 16.968 0.002 0.038 0.038 0.955 0.271
21 21.218 0.003 0.040 0.040 0.952 0.280
25 18.707 -0.024 0.038 0.045 0.957 0.251
30 16.976 0.001 0.037 0.037 0.957 0.279
35 24.519 -0.005 0.034 0.034 0.939 0.233
40 33.199 -0.014 0.036 0.039 0.952 0.243
Uniform 0.951 0.407
rd2d.dist kink = "on" 1 15.007 0.009 0.075 0.075 0.954 0.674
5 11.715 0.004 0.072 0.072 0.933 0.611
10 8.572 -0.002 0.075 0.075 0.935 0.655
15 7.434 -0.002 0.080 0.080 0.923 0.691
21 9.296 0.004 0.087 0.087 0.924 0.738
25 8.196 -0.002 0.074 0.074 0.928 0.645
30 7.437 0.001 0.081 0.081 0.947 0.714
35 10.742 0.002 0.071 0.071 0.945 0.593
40 14.545 -0.001 0.074 0.074 0.950 0.632
Uniform 0.877 1.071
@@
rd2d 1 17.329 0.006 0.048 0.048 0.953 0.289
5 15.394 0.006 0.036 0.037 0.964 0.220
10 13.339 0.004 0.034 0.034 0.943 0.197
15 13.789 0.004 0.030 0.030 0.953 0.177
21 15.194 -0.010 0.045 0.046 0.952 0.297
25 13.954 -0.006 0.028 0.029 0.958 0.164
30 13.086 -0.003 0.033 0.033 0.950 0.192
35 14.511 -0.003 0.036 0.036 0.950 0.206
40 16.851 -0.004 0.046 0.046 0.943 0.264
Uniform 0.953 0.315
rd2d.dist kink = "off" 1 34.144 0.039 0.038 0.055 0.927 0.259
5 24.612 0.016 0.039 0.042 0.942 0.260
10 18.392 0.002 0.037 0.037 0.948 0.271
15 14.216 -0.002 0.042 0.042 0.957 0.320
21 21.137 -0.005 0.038 0.039 0.956 0.282
25 16.227 -0.025 0.045 0.052 0.950 0.288
30 13.810 -0.004 0.047 0.047 0.944 0.337
35 19.163 -0.019 0.045 0.048 0.927 0.294
40 31.515 -0.070 0.041 0.082 0.946 0.259
Uniform 0.948 0.460
rd2d.dist kink = "on" 1 14.959 0.017 0.081 0.083 0.942 0.677
5 10.783 0.001 0.079 0.079 0.940 0.663
10 8.058 0.003 0.082 0.082 0.942 0.691
15 6.228 -0.001 0.095 0.095 0.919 0.821
21 9.260 -0.001 0.085 0.085 0.924 0.749
25 7.109 -0.000 0.088 0.088 0.917 0.734
30 6.050 -0.003 0.104 0.104 0.916 0.840
35 8.395 -0.003 0.089 0.089 0.924 0.754
40 13.807 -0.017 0.078 0.079 0.943 0.664
Uniform 0.846 1.197

### v2

@@ 3.6 Simulation Evidence
rd2d 1 15.764 0.005 0.053 0.054 0.953 0.313
5 13.885 0.003 0.042 0.042 0.958 0.243
10 11.997 0.002 0.038 0.038 0.952 0.218
15 12.706 0.001 0.033 0.033 0.949 0.191
21 13.788 -0.003 0.049 0.049 0.957 0.329
25 12.908 0.001 0.031 0.031 0.950 0.177
30 11.688 -0.001 0.036 0.036 0.948 0.212
35 13.012 0.001 0.039 0.039 0.957 0.228
40 15.255 0.001 0.049 0.049 0.960 0.289
Uniform 0.948 0.348
rd2d.dist kink = "off" 1 34.136 0.031 0.038 0.049 0.940 0.261
5 26.747 0.016 0.035 0.038 0.950 0.240
10 19.478 0.003 0.035 0.035 0.953 0.258
15 16.905 0.001 0.038 0.038 0.957 0.275
21 21.151 0.002 0.038 0.038 0.960 0.283
25 18.524 -0.021 0.039 0.044 0.955 0.255
30 16.986 -0.000 0.037 0.037 0.948 0.280
35 24.532 -0.005 0.033 0.033 0.956 0.233
40 33.070 -0.015 0.035 0.038 0.953 0.244
Uniform 0.946 0.408
rd2d.dist kink = "on" 1 14.955 0.012 0.079 0.080 0.954 0.729
5 11.718 0.003 0.072 0.072 0.948 0.646
10 8.533 0.001 0.078 0.078 0.954 0.698
15 7.406 -0.002 0.083 0.083 0.948 0.753
21 9.266 -0.003 0.083 0.083 0.944 0.810
25 8.116 0.005 0.076 0.077 0.955 0.697
30 7.442 -0.001 0.079 0.079 0.951 0.761
35 10.748 0.002 0.067 0.067 0.948 0.626
40 14.488 -0.001 0.075 0.075 0.948 0.667
Uniform 0.89 1.073
@@
rd2d 1 15.730 0.004 0.052 0.053 0.958 0.315
5 13.803 0.005 0.041 0.041 0.965 0.244
10 11.954 0.003 0.038 0.038 0.943 0.218
15 12.600 0.001 0.033 0.033 0.944 0.193
21 13.716 -0.008 0.049 0.049 0.962 0.331
25 12.833 -0.003 0.031 0.031 0.952 0.177
30 11.679 0.000 0.035 0.034 0.962 0.212
35 12.966 -0.001 0.041 0.041 0.954 0.230
40 15.312 -0.007 0.050 0.050 0.942 0.288
Uniform 0.942 0.350
rd2d.dist kink = "off" 1 34.303 0.040 0.038 0.055 0.927 0.260
5 24.471 0.016 0.037 0.041 0.960 0.262
10 18.392 -0.001 0.036 0.036 0.951 0.273
15 14.179 -0.003 0.042 0.042 0.944 0.324
21 21.170 -0.004 0.037 0.038 0.959 0.283
25 16.036 -0.020 0.045 0.049 0.943 0.293
30 13.771 -0.002 0.043 0.043 0.947 0.341
35 19.017 -0.016 0.043 0.046 0.954 0.297
40 31.541 -0.071 0.042 0.083 0.935 0.260
Uniform 0.949 0.461
rd2d.dist kink = "on" 1 15.028 0.015 0.079 0.080 0.950 0.718
5 10.721 0.005 0.074 0.074 0.958 0.710
10 8.058 0.000 0.082 0.082 0.937 0.746
15 6.212 -0.008 0.097 0.097 0.942 0.903
21 9.275 -0.001 0.084 0.084 0.945 0.811
25 7.026 0.001 0.088 0.088 0.950 0.810
30 6.033 0.006 0.102 0.102 0.944 0.947
35 8.331 -0.001 0.090 0.090 0.957 0.817
40 13.818 -0.020 0.079 0.081 0.952 0.708
Uniform 0.822 1.199
@@ 4 Conclusion
5 Acknowledgments
We thank Xinwei Ma and Filippo Palomba for insightful discussions. Cattaneo and Titiunik gratefully acknowledge financial support from the National Science Foundation (SES-2019432 and SES-2241575). Cattaneo gratefully acknowledge financial support from the Data-Driven Social Science initiative at Princeton University.